OpenStudy (anonymous):

How do I solve for x in the equation y=-x+4?

pooja195 (pooja195):

Ok so you said y-4=x

OpenStudy (anonymous):

No, y-4 = -x

pooja195 (pooja195):

^right divide everything by -1 what are you left with?

OpenStudy (anonymous):

y+4 = x so x= y+4 right ?

pooja195 (pooja195):

Well the y was orginally positive so y/-1= ?

OpenStudy (anonymous):

I have to do substitution The page says to solve each system by substitution The problem is y=-x+4 and then y=3x

OpenStudy (anonymous):

I have to solve for x and then substitute it into the other equation.

pooja195 (pooja195):

ooohhh that makes sense y=-x+4 <---equation 1 y=3x <---equation 2 We know the value of y take equation 1 and plug it into equation 2 \[\huge~\rm~\bf~\color{blue}{ x+4}=3x\] Solve for x

pooja195 (pooja195):

Are you stuck?

pooja195 (pooja195):

Start by subtracting x from both sides \[\huge~\rm~ x- x+4=3x-x\]

pooja195 (pooja195):

\[\huge~\rm~ 4=3x-x\] 3x-x=?

OpenStudy (anonymous):

its y = -x+4 not x+4

pooja195 (pooja195):

ooohhh that makes sense y=-x-4 <---equation 1 y=3x <---equation 2 We know the value of y take equation 1 and plug it into equation 2 \[\huge~\rm~\bf~\color{blue}{ x-4}=3x\] Solve for x Like this?